OpenStudy (anonymous):

I will fan and medal! A game is played with the following rules. Two ordinary number cubes, each with six sides labeled 1 to 6, are rolled. After each roll a player is given a certain number of points depending on the outcome. If the sum is 2 or 12, the player receives +6 points If the sum is 3 or 11, the player receives +2 points. If the sum is any other number, the player receives –1 point. What is the expected value of the number of points a player will receive after a roll?

OpenStudy (anonymous):

Do you want someone to help you find the answer, or would you just like the answer?

OpenStudy (anonymous):

Having 36 possible outcomes; 2 = 6 points 4 = 2 points 30 = -1 point The expected number of points received is (2*6 + 4*2 + 30*-1)/36 = -10/36 = -5/18